OpenStudy (anonymous):

Evaluate the iterated integral. 2 2 ∫ ∫ xydydx 1 x Explain please.

sam (.sam.):

\[\int\limits_{1}^{2}\int\limits_{x}^{2}xy~dydx\] Solve with respect to y, treating x as a constant \[\huge \int\limits\limits_{1}^{2}[\frac{xy^{2}}{2}]_{x}^{2} dx\] \[\huge \int\limits\limits\limits_{1}^{2}2x-\frac{x^3}{2}dx\] \[[\huge x^2-\frac{x^4}{(4)2}]_{1}^{2}\] \[\huge =\frac{9}{8}\]

sam (.sam.):

where you don't understand just say it out :)

OpenStudy (anonymous):

how did you know to begin solving with respect to y instead of x?

sam (.sam.):

the 'dy' tells

OpenStudy (anonymous):

thank you so much i understand it =P!

sam (.sam.):

Its basically separated like this \[\huge (\int\limits_{1}^{2}[\int\limits_{x}^{2}xydy]dx)\]

sam (.sam.):

do the inner part first

OpenStudy (anonymous):

for the second part when you pluged in 2 and x did you plug it into the y or the x?

sam (.sam.):

the 'y' because of 'dy'

OpenStudy (anonymous):

thank you i completed the problem! =]